OpenStudy (anonymous):

how do i solve this by factoring? w(w-3)=108

OpenStudy (inkyvoyd):

w^2-3w-108=0 Now we are looking for the factors of 108 that are 3 from each other.

OpenStudy (inkyvoyd):

Let's try 12*9

OpenStudy (anonymous):

w^2-3w-108=0 w^2-12w+9w-108=0 W(w-12)+9(w-12)=0(w-12)(w-9)=0 w=12,9

OpenStudy (inkyvoyd):

(w+9)(w-12) or (w-9)(w+12)

OpenStudy (inkyvoyd):

since we see -3w, we have (w+9)(w-12).

OpenStudy (inkyvoyd):

Not what @Somjit said xD.

OpenStudy (anonymous):

k got it, thx sooo much